OpenStudy (anonymous):

factor...30s^2+19s-63

OpenStudy (anonymous):

a*x^2+b*x+c=0 D = b^2-4*ac......=89^2 1x2=-b+-sqrt(D)/2a

OpenStudy (anonymous):

so the result is 7/6 and 9/5